How do I find the standard form of a circle?
The question states
x^2 + Y^2 +6x - 6y - 7 = 0

Answer by fractalier(6550)

Let us complete the square twice, once in x and once in y and get
x^2 + y^2 + 6x - 6y - 7 = 0
x^2 + 6x + 9 + y^2 - 6y + 9 - 7 = 0 + 9 + 9
(x + 3)^2 + (y - 3)^2 = 25 = 5^2
Center at (-3, 3) and radius 5.